OpenStudy (anonymous):

through (-2,0) perpendicular to 8x-3y=7 (a) I need it in standard form (b) in slope interept form

OpenStudy (anonymous):

the orginal line has slope 8/3, so we will use -3/8. y=-3/8(x+2), so: a) 3x+8y=-6 b) y=-3/8x-3/4

OpenStudy (anonymous):

THANK YOU SO MUCH FOR YOUR HELP!!!!